1) 42 is 40% of what number?
2) 15 is what percent of 3?
3) The sum of the page numbers on facing pages of a book is 157; what are the page numbers?

Answer by checkley77(12569)

.42*40=16.8 answer
15/3=5 or 500% answer.
x+x+1=157
2x=157-1
2x=156
x=156/2
x=78 answer.